# An AI System Evaluation Framework for Advancing AI Safety: Terminology, Taxonomy, Lifecycle Mapping

### v2

@@ Abstract.
copyright: rightsretaineddoi: 10.1145/3664646.3664766journalyear: 2024submissionid: fsews24aiwaremain-p44-pisbn: 979-8-4007-0685-1/24/07conference: Proceedings of the 1st ACM International Conference on AI-Powered Software; July 15–16, 2024; Porto de Galinhas, Brazilbooktitle: Proceedings of the 1st ACM International Conference on AI-Powered Software (AIware ’24), July 15–16, 2024, Porto de Galinhas, Brazil
@@ 2. Harmonised Terminology
Benchmarking, a crucial evaluation type, is especially pivotal for evaluating General AI which spans various tasks. This breadth necessitates a comprehensive examination of aspects such as quality/risk (e.g., fairness and bias), accuracy/correctness (e.g., a translation AI’s fidelity and the translation software’s operational integrity), and (intended and unintended) capabilities (Weidinger et al., 2023). By systematically comparing against various baselines and rivals, benchmarking ensures not only performance excellence but also ethical integrity.
@@ 3. AI System Evaluation: A Taxonomy
Fig. 1 presents a taxonomy for evaluating AI systems at component and system levels, emphasising key areas and highlighting differences between model versus system and Narrow AI versus General AI.
Figure 1. A Taxonomy for AI System Evaluation (Model vs. System Perspective)
@@ 3.1. Component-Level Evaluation
Evaluation at this level includes both AI (e.g., AI model) and non-AI components.
@@ 3.1.3. AI Components Evaluation - Model
Figure 2. AI System Evaluation (Supply Chain Perspective)
@@ 3.1.4. Safety Guardrails Evaluation
The evaluation addresses three critical attributes: Generalisability across different AI models/systems and contexts, ensuring they remain effective under varied operational scenarios. Customisability is also crucial, allowing guardrails to be “fine-tuned” to the specific needs of each deployment (Zhang et al., 2024). For example, autonomous driving guardrails might issue a warning in certain region when hands are off the wheel but initiate slowing down in another, reflecting deployment contexts. Lastly, standards alignment guarantees that guardrails adhere to established safety, ethical, and regulatory frameworks, making them not just technically sound but also ethically responsible and compliant with legal requirements.

### v3

@@ 2. Harmonised Terminology
Benchmarking, a crucial evaluation type, is especially pivotal for evaluating General AI which spans various tasks. This breadth necessitates a comprehensive examination of aspects such as quality/risk (e.g., fairness and bias), accuracy/correctness (e.g., a translation AI’s fidelity and the translation software’s operational integrity), and (intended and unintended) capabilities (Weidinger et al., 2023). By systematically comparing against diverse baselines and competitors, benchmarking ensures not only performance excellence but also ethical integrity across different contexts (e.g., scenarios, use cases). This context-based evaluation is essential for obtaining meaningful and actionable insights, ensuring the relevance and applicability of the benchmarking results to real-world deployments.
@@ 3. AI System Evaluation: A Taxonomy
Fig. 1 presents a taxonomy for evaluating AI systems at component and system levels, emphasising key areas.
Figure 1. A Taxonomy for AI System Evaluation
@@ 3.1. Component-Level Evaluation
Evaluation at this level includes both AI and non-AI components.
@@ 3.1.3. AI Components Evaluation - Model
Figure 2. AI System Evaluation (Supply Chain Perspective)
@@ 3.1.4. Safety Guardrails Evaluation
The evaluation addresses four critical attributes: Generalisability across different AI models/systems and contexts, ensuring they remain effective under varied operational scenarios; Customisability, which allows guardrails to be “fine-tuned” to the specific needs of each deployment (Zhang et al., 2024). For example, autonomous driving guardrails might issue a warning in a certain region when hands are off the wheel but initiate slowing down in another, reflecting deployment contexts; Interpretability, which determines whether human users can understand the guardrails. This is crucial for ensuring that users can comprehend the safety measures in place and trust the system’s operations. Clear and understandable guardrails enhance user trust and facilitate better interaction with the AI system; and Standards alignment, ensuring that guardrails adhere to established safety, ethical, and regulatory frameworks, making them not just technically sound but also ethically responsible and compliant with legal requirements.